# Buyer-Initiated Auction Mechanism for Data Redemption in Machine Unlearning

### v3

@@ Abstract
The rapid growth of artificial intelligence (AI) has raised privacy concerns over user data, leading to regulations like the General Data Protection Regulation (GDPR) and the California Consumer Privacy Act (CCPA). With the essential toolbox provided by machine unlearning, AI service providers are now able to remove user data from their trained models as well as the training datasets, so as to comply with such regulations. However, extensive data redemption can be costly and degrade model accuracy. To balance the cost of unlearning and the privacy protection, we propose a buyer-initiated auction mechanism for data redemption, enabling the service provider to purchase data from willing users with appropriate compensation. This approach does not require the server to have any a priori knowledge about the users’ privacy preference, and provides an efficient solution for maximizing the social welfare in the investigated problem.
Index Terms:
Data privacy, data redemption, machine unlearning, auction mechanism.
@@ I Introduction
The rapid development of artificial intelligence (AI) and machine learning (ML) technologies over the past decade has brought about a prospery of intelligent applications and a new era of digital economy, recently supplemented by the explosive blowout of large language models (LLMs). As the feul of this new techno-economic paradigm, high-quality data has become a valuable asset and in thirst demand. In practice, the data used for training ML models is often collected from the users of the applications. However, the data collection process may raise privacy concerns, as the data may contain sensitive information about the users, and the models therewith trained may leak or abuse such information against the users’ will.
@@
To enable removing user’s data from the server in both senses of the training dataset and the trained model, the technologies of machine unlearning have been developed and attracted much attention [4]. Despite the technical success, however, an excessive unlearning of user data can be not only computationally costly for the AI service provider, but also significantly degrading the model’s accuracy, which also harm the users’ quality of experience [5]. Therefore, a rigid unlearning policy may be the long run lead against both interests of the service provider and the users.
@@
In this work, we propose to invoke buyer-initiated ascending auction mechanism to address the above limitations. In a multi-round auction, the server gradually raises the unit price for purchasing data, and each user independently decides how much data it is willing to sell at the offered price. After the users claiming, the server determines how much data it is willing to buy at the current price, before announcing the new price for the next round or closing the auction. This auction mechanism enables a flexible and progressive pricing to convey the non-linear cost/utility functions of both the server and the users, and requires no prior knowledge on the server side about the distribution of users’ privacy parameters.
The remainding part of this paper is organized as follows. In Section II, we present the system model, including the cost and utility functions of the server and the users, respectively. In Section III, we analyze the server’s incentive to purchase data and the users’ incentive to sell data in an arbitrary auction round. In Section IV, we propose the buyer-initiated auction mechanism, and discuss the optimal decision strategies for the server and the users. Subsequently, we numerically evaluate our proposed approach with simulation campaigns in Sec. V, before finally concluding the paper in Section VI.
@@ II-A Server Cost Model
Considering the cost model in [9] that $C(x)=\alpha A(x)+\beta T(x)$, where $x=\sum\limits_{i=i}^{I}x_{i}$, $x_{i}$ is the amount of data redemption for user $i\in\mathcal{I}$, $A(x)$ is the accuracy degradation of the model caused by unlearning, and $T(x)$ the computing time for executing the unlearning task. Note that $\alpha,A,\beta,T$ are all non-negative. Defining the total data amount for each user $i$ as $d_{i}$, and their sum $d=\sum\limits_{i=1}^{I}d_{i}$, according to [9]:
@@
$T(x)$ =$=\begin{cases}0&\text{if }x=0,\\ T_{0}(d-x)&\text{if }x\in(0,d].\end{cases}$ (2)
From the server’s perspective it is more convenient to focus on the amount of data to keep, i.e. to buy from the users, rather than the amount of data to redeem. So we define that $y_{i}=d_{i}-x_{i}$ and $y=d-x$, and rewrite the cost function as
@@
$T(y)$ =$=\begin{cases}T_{0}y&\text{if }y\in[0,d),\\ 0&\text{if }y=d.\end{cases}$ (5)
@@ III-B User’s Incentive to Sell Data
IV Multi-Round Auction Mechanism
IV-A Buyer-Initiated English Auction
To balance the server’s cost of unlearning and the users’ privacy utility, we propose a buyer-initiated ascending auction mechanism, which can be briefly summarized as follows:
@@ IV-A Buyer-Initiated English Auction
    At the beginning of each round $t$, the server updates its price $B^{t}$ for buying unit data from the users. We focus in this study on the ascending auction format, where the data price is increasing w.r.t. time, i.e., for all $t<\tau$, $B^{t}<B^{\tau}$. It also estimates its demand for data $\eta^{t}=\eta(y^{t},B^{t})$, which is unknown to the users. If the demand drops to zero, the auction is terminated. Otherwise, the price is announced to all users.
@@
    The server purchases the data from all users who are willing to sell at the announced price. If the total supply exceeds its demand, it will only purchase its demanded amount of data. Otherwise, it will buy all data supplied.
@@
    If the server has collected the target amount of data $y_{\mathrm{max}}$, the auction is terminated. Otherwise, $t\leftarrow t+1$, and go back to step 1) for the next round.
For the sake of low communication overhead and privacy protection, our approach is to let the users make decisions independently from each other. Thus, each user $i$ is considered blind in the sense that at each period $t$, $i$ only knows the current data price $B^{t}$, its own privacy parameter $\lambda_{i}$, its own data amount $d_{i}$, and its own history of data selling $(y_{i}^{\tau})_{\tau<t}$. As we are considering an ascending auction, the price announced by the server along the time assembles an ascending series $(B^{\tau})_{1\leqslant\tau<\infty}$.
@@ IV-B Users’ Selling Strategy
Noticing the $t$-monotonically increasing price $B^{t}$, when a user has incentive to sell data in round $t$, it has also the alternative option of keeping the data to sell it at higher price in a future round. Herewith we discuss the optimal selling strategy of users in such multi-round auction.
Suppose the auction is now at the beginning of round $t$. Each user $i$’s current remaining data that can be sold is $x_{i}^{t}\geqslant 0$. Given the information that $i$ knows, $i$ forms a belief $\mu_{i}^{t}\in[0,1]$, which represents a subjective probability that the auction is terminated after round $t$. Thus, $\mu_{i}^{t}$ is a function about server’s remaining demand $\eta^{t}$ at $t$, and the supply $q_{j}^{t}$ of all users $j\in\mathcal{I}$.
@@ IV-C Handling the Oversupply
Due to the wide-sense $B$-monotonicity of $\eta$ (decreasing) and $q_{i}$ (increasing) for all $i\in\mathcal{I}$, when the announced price $B^{t}$ is sufficiently high, it may occur the case of $\eta^{t}<\sum\limits_{i=1}^{I}q_{i}^{t}$, i.e. the total supply from users exceeds the server’s demand. In this case, the server shall allocate its demand $\eta$ to the different selling users.
@@
Remark that the oversupply will certainly trigger the satisfaction of server’s optimal amount of data to keep $y_{\mathrm{max}}$, so it will only occur once, i.e., in the last round of the auction.
@@ IV-D Handling the Discontinuity in Cost
When $y_{\mathrm{max}}<d$, due to this jump discountinuity, the actual server’s cost of unlearning all user data of amount $d$ is $C(d)=\alpha A(d)$, which can be significantly lower than $C(y_{\mathrm{max}})$. Thus, after achieving $y^{t}=y_{\mathrm{max}}$, the server may still have an incentive to see if it worth buying all the remaining data from the users. Nevertheless, since the server does not have full knowledge about every user’s privacy parameter, it cannot simply estimate the optimal price to offer. It is neither rational to simply continue purchasing more data at prices higher than the marginal cost of unlearning. To address this issue, we propose a post-auction procedure, where after buying $y_{\mathrm{max}}$ amount of data, the server will still keep announcing new prices to the users like if the auction continues, but not purchasing. Instead, it observes if all users are willing to sell all remaining data at the new price. Only if so, the server will purchase all remaining data at the last announced price, and terminates the post-auction step. Otherwise, no purchase will be made even if the supply is non-zero, and the server simply updates the price. This procedure continues until the price $B^{t}$ is raised so high that the cost to buy all remaining data exceeds the unlearning cost saved therewith, i.e. when $B^{t}\cdot\left(d-y_{\mathrm{max}}\right)>C\left(y_{\mathrm{max}}\right)-C(d)$.
@@ IV-E Proposed Algorithm
Summarizing the above discussions, we propose the algorithm for the buyer-initiated ascending auction in Algorithm 1. Especially, for practical implementation in real use scanrios, we consider discrete user data sets with unit size of $\Delta d$, so that all trades of user data can only be executed by amount of integer times of $\Delta d$.
1 Input: $a,A_{1},A_{2},A_{3},T_{0},\alpha,\beta,B^{0},\Delta B,[d_{1},d_{2},\dots,d_{I}],\Delta d$,
@@
3 while $\eta^{t}>0$ do // Auction phase
@@ 15       $t\leftarrow t+1$
18while $B^{t}\cdot\left(d-y^{t}\right)\leqslant C\left(y^{t}\right)-C(d)$ do // Post-auction phase
@@ 25       $t\leftarrow t+1$
Algorithm 1 Buyer-Initiated English Auction
@@ V-A Simulation Setup
$\Delta d$ $1$ Unit data amount
@@
System $N$ $1000$ Monte Carlo runs
$[a,A_{1},A_{2},A_{3}]$ $[e,0.1,3.33\times 10^{-5},0]$ Accuracy parameters
@@
Cost $[\alpha,\beta]$ [1500,1] Weight factors
@@ V-B Oversupply Handling Strategies
The simulation results are shown in Table II. As we have expected, the selection of oversupply-handling strategy does not affect the server’s payoff at all (difference within $0.2\%$), while the minor-seller-first strategy outperforms the other three regarding the users’ payoff and social welfare by a very slight superiority (within $2\%$). This is not suprising as the oversupply handling is triggered only in the last auction round, which takes only a small portion of the total auction procedure.
@@ V-C Baseline Methods
To comparatively benchmark the performance of our proposed auction approach against state-of-the-art solutions, we also implemented three baseline methods:
@@
    Bi-stage pricing (BSP): This is the solution proposed in [9], where the server solves an optimal universal price and announce it to all users, each user then independently decides how much data to sell at that price.
@@ V-D Benchmark Test
We conducted a benchmark test to compare the performance of our proposed auction approach with the aforementioned baseline methods. Specifically, the auction approach was configured to handle oversupply with the minor-sellers-first strategy. We tested all methods at different informed ratios from $0\%$ to $100\%$ with $10\%$ step length, and measured the server’s payoff, users’ payoff, and social welfare. The results of Monte Carlo tests are shown in Fig. 2.
@@
As we can see, our proposed auction approach outperforms all the other methods in terms of social welfare. While the social welfare brought by the rigid GDPR solution drops at high informed ratio due to the high unlearning cost, both our auction approach and the BSP solution exhibit an social welfare steadily increasing along with the informed ratio. Particularly, in comparison to BSP, our method significantly improves the server’s payoff at the cost of a less reduction in users’ payoff, leading to its superiority in the overall social welfare. This shall be explained in two aspects:
@@
    The multi-round auction mechanism allows the server to explore the users’ non-linear privacy function over multiple intervals of data trading amount, which provides a closer approximation of the privacy utility loss by the payment. Compared to the BSP solution that provides only one optimum average price, the auction mechanism is more cost-profitable for the buyer (server), but less profitable for the users.
@@ VI Conclusions
In this paper, we have proposed a buyer-initiated ascending auction framework for data redemption in machine unlearning. We have analyzed the incentives of buying and selling data of the server and users, respectively, and derived the optimal selling strategy of users. Compared with state-of-the-art baselines, our proposed solution does not require the server to possess any knowledge about the users’ privacy parameter, and yet achieving a higher social welfare, which is demonstrated by numerical simulations.

### v4

# Quotation-Based Data Retention Mechanism for Data Privacy in LLM-Empowered Network Services

@@ Abstract
The deployment of large language models (LLMs) for next-generation network optimization introduces novel data governance challenges. mobile network operators (MNOs) increasingly leverage generative artificial intelligence (AI) for traffic prediction, anomaly detection, and service personalization, requiring access to users’ sensitive network usage data–including mobility patterns, traffic types, and location histories. Under the General Data Protection Regulation (GDPR), the California Consumer Privacy Act (CCPA), and similar regulations, users retain the right to withdraw consent and demand data deletion. However, extensive machine unlearning degrades model accuracy and incurs substantial computational costs, ultimately harming network performance for all users. We propose an iterative price discovery mechanism enabling MNOs to compensate users for data retention through sequential price quotations. The server progressively raises the unit price for retaining data while users independently determine their supply at each quoted price. This approach requires no prior knowledge of users’ privacy preferences and efficiently maximizes social welfare across the network ecosystem.
@@ I Introduction
The integration of large language models (LLMs) and generative artificial intelligence (AI) into next-generation mobile networks has emerged as a key enabler for intelligent network management and operations [1]. mobile network operators (MNOs) deploy LLMs for predictive maintenance, automated troubleshooting, traffic forecasting, and personalized service delivery. These applications require training on extensive user data–including cell handover records, packet inspection logs, geographic mobility traces, and application usage patterns–to achieve acceptable accuracy.
However, the data collection process may raise privacy concerns, as the data may contain sensitive information about the users, and the models therewith trained may leak or abuse such information against the users’ will. In network operations contexts, this tension intensifies: user data fed into LLM training directly influences quality of service (QoS) for the entire user base, creating interdependencies between individual privacy and collective network performance.
@@
To enable removing user data from LLM training pipelines–both from stored datasets and deployed models–MNOs must implement machine unlearning [5]. Despite technical feasibility, excessive unlearning creates dual challenges for network operators: (1) computational overhead in retraining resource- intensive LLMs, and (2) degraded predictive accuracy that cascades into worse traffic management, increased latency, and higher outage rates [6]. A rigid unlearning policy paradoxically harms both MNO operational efficiency and user experience, necessitating market-based alternatives.
@@
In this work, we propose an iterative price discovery mechanism based on sequential quotations to address the above limitations. The server (MNO) progressively quotes higher unit prices for data retention, and users independently determine their willingness to supply data at each price point. This quotation-based approach enables a flexible and progressive pricing to convey the non-linear cost/utility functions of both the server and the users, and requires no prior knowledge on the server side about the distribution of users’ privacy parameters.
The remainding part of this paper is organized as follows. In Section II, we present the system model, including the cost and utility functions of the server and the users, respectively. In Section III, we analyze the server’s incentive to purchase data and the users’ incentive to sell data in an arbitrary quotation round. In Section IV, we propose the iterative quotation-based price discovery mechanism, and discuss the optimal decision strategies for the server and the users. Subsequently, we numerically evaluate our proposed approach with simulation campaigns in Sec. V, before finally concluding the paper in Section VI.
@@ II System Model
We model the interaction between an MNO deploying LLMs for network optimization and its user base. The MNO (henceforth ”server”) operates an LLM trained on aggregated user data to perform tasks such as:
  • Traffic prediction: Forecasting bandwidth demand for proactive resource allocation;
  • Anomaly detection: Identifying network faults or security threats;
  • Service personalization: Recommending optimal connectivity plans.
Each user $i\in\mathcal{I}$ has contributed data di (measured in normalized units, e.g., number of mobility records or session logs). Under GDPR Article 17, users may request data deletion, triggering machine unlearning obligations. The following cost and utility models capture the economic tradeoffs.
@@ II-A Server Cost Model
Considering the cost model in [6] that $C(x)=\alpha A(x)+\beta T(x)$, where $x=\sum\limits_{i=1}^{I}x_{i}$, $x_{i}$ is the amount of data redemption for user $i\in\mathcal{I}$, $A(x)$ is the accuracy degradation of the model caused by unlearning, and $T(x)$ the computing time for executing the unlearning task. Note that $\alpha,A,\beta,T$ are all non-negative. Defining the total data amount for each user $i$ as $d_{i}$, and their sum $d=\sum\limits_{i=1}^{I}d_{i}$, according to [6]:
@@
$T(x)$ $=\begin{cases}0&\text{if }x=0,\\ T_{0}(d-x)&\text{if }x\in(0,d].\end{cases}$ (2)
From the server’s perspective it is more convenient to focus on the amount of data to keep, i.e. to buy from the users, rather than the amount of data to redeem. So we define that $y_{i}=d_{i}-x_{i}$ and y=dxy=d-x, and rewrite the cost function as
@@
$T(y)$ $=\begin{cases}T_{0}y&\text{if }y\in[0,d),\\ 0&\text{if }y=d.\end{cases}$ (5)
@@ III-B User’s Incentive to Sell Data
IV Quotation-Based Price Discovery Mechanism
IV-A Quotation-Based Price Discovery Protocol
To balance the server’s cost of unlearning and the users’ privacy utility, we propose an iterative price discovery mechanism through sequential quotations. The protocol operates as follows:
@@ IV-A Quotation-Based Price Discovery Protocol
    At the beginning of each round $t$, the server updates its price $B^{t}$ for buying unit data from the users. We focus in this study on the ascending quotation format, where the data price is increasing w.r.t. time, i.e., for all $t<\tau$, $B^{t}<B^{\tau}$. It also estimates its demand for data $\eta^{t}=\eta(y^{t},B^{t})$, which is unknown to the users. If the demand drops to zero, the quotation is terminated. Otherwise, the price is quoted to all users.
@@
    The server purchases the data from all users who are willing to sell at the quoted price. If the total supply exceeds its demand, it will only purchase its demanded amount of data. Otherwise, it will buy all data supplied.
@@
    If the server has collected the target amount of data $y_{\mathrm{max}}$, the quotation is terminated. Otherwise, $t\leftarrow t+1$, and go back to step 1) for the next round.
For the sake of low communication overhead and privacy protection, our approach is to let the users make decisions independently from each other. Thus, each user $i$ is considered blind in the sense that at each period $t$, $i$ only knows the current data price $B^{t}$, its own privacy parameter $\lambda_{i}$, its own data amount $d_{i}$, and its own history of data selling $(y_{i}^{\tau})_{\tau<t}$. As we are considering an ascending quotation, the price quoted by the server along the time assembles an ascending series $(B^{\tau})_{1\leqslant\tau<\infty}$.
@@ IV-B Users’ Selling Strategy
Noticing the $t$-monotonically increasing price $B^{t}$, when a user has incentive to sell data in round $t$, it has also the alternative option of keeping the data to sell it at higher price in a future round. Herewith we discuss the optimal selling strategy of users in such multi-round quotation.
Suppose the quotation is now at the beginning of round $t$. Each user $i$’s current remaining data that can be sold is $x_{i}^{t}\geqslant 0$. Given the information that $i$ knows, $i$ forms a belief $\mu_{i}^{t}\in[0,1]$, which represents a subjective probability that the quotation is terminated after round $t$. Thus, $\mu_{i}^{t}$ is a function about server’s remaining demand $\eta^{t}$ at $t$, and the supply $q_{j}^{t}$ of all users $j\in\mathcal{I}$.
@@ IV-C Handling the Oversupply
Due to the wide-sense $B$-monotonicity of $\eta$ (decreasing) and $q_{i}$ (increasing) for all $i\in\mathcal{I}$, when the quoted price $B^{t}$ is sufficiently high, it may occur the case of $\eta^{t}<\sum\limits_{i=1}^{I}q_{i}^{t}$, i.e. the total supply from users exceeds the server’s demand. In this case, the server shall allocate its demand $\eta$ to the different selling users.
@@
Remark that the oversupply will certainly trigger the satisfaction of server’s optimal amount of data to keep $y_{\mathrm{max}}$, so it will only occur once, i.e., in the last round of the quotation.
@@ IV-D Handling the Discontinuity in Cost
When $y_{\mathrm{max}}<d$, due to this jump discountinuity, the actual server’s cost of unlearning all user data of amount $d$ is $C(d)=\alpha A(d)$, which can be significantly lower than $C(y_{\mathrm{max}})$. Thus, after achieving $y^{t}=y_{\mathrm{max}}$, the server may still have an incentive to see if it worth buying all the remaining data from the users. Nevertheless, since the server does not have full knowledge about every user’s privacy parameter, it cannot simply estimate the optimal price to offer. It is neither rational to simply continue purchasing more data at prices higher than the marginal cost of unlearning. To address this issue, we propose a post-quotation procedure, where after buying $y_{\mathrm{max}}$ amount of data, the server will still keep announcing new prices to the users like if the quotation continues, but not purchasing. Instead, it observes if all users are willing to sell all remaining data at the new price. Only if so, the server will purchase all remaining data at the last quoted price, and terminates the post-quotation step. Otherwise, no purchase will be made even if the supply is non-zero, and the server simply updates the price. This procedure continues until the price $B^{t}$ is raised so high that the cost to buy all remaining data exceeds the unlearning cost saved therewith, i.e. when $B^{t}\cdot\left(d-y_{\mathrm{max}}\right)>C\left(y_{\mathrm{max}}\right)-C(d)$.
@@ IV-E Proposed Algorithm
Summarizing the above discussions, we propose an iterative quotation-based price discovery protocol in Algorithm 1. Especially, for practical implementation in real use scanrios, we consider discrete user data sets with unit size of $\Delta d$, so that all trades of user data can only be executed by amount of integer times of $\Delta d$.
1Input: $a,A_{1},A_{2},A_{3},T_{0},\alpha,\beta,B^{0},\Delta B,[d_{1},d_{2},\dots,d_{I}],\Delta d$,
@@
3 while $\eta^{t}>0$ do // Quotation phase
@@ 15  $t\leftarrow t+1$
18while $B^{t}\cdot\left(d-y^{t}\right)\leqslant C\left(y^{t}\right)-C(d)$ do // Post-quotation phase
@@ 25  $t\leftarrow t+1$
Algorithm 1 Iterative Price Discovery Protocol
@@ II $10$ Number of users
$\Delta d$ 11 Unit data amount
@@
System NN $1000$ Monte Carlo runs
$a,A_{1},A_{2},A_{3}] [$$e,0.1,3.33\times 10^{-5},0]$ Accuracy parameters
@@
Cost $\alpha,\beta] [$1500,1] Weight factors
@@ V-B Oversupply Handling Strategies
The simulation results are shown in Table II. As we have expected, the selection of oversupply-handling strategy does not affect the server’s payoff at all (difference within $0.2\%$), while the minor-seller-first strategy outperforms the other three regarding the users’ payoff and social welfare by a very slight superiority (within $2\%$). This is not suprising as the oversupply handling is triggered only in the last quotation round, which takes only a small portion of the total quotation procedure.
@@ V-C Baseline Methods
To comparatively benchmark the performance of our proposed quotation approach against state-of-the-art solutions, we also implemented three baseline methods:
@@
    Bi-stage pricing (BSP): This is the solution proposed in [6], where the server solves an optimal universal price and quotes it to all users, each user then independently decides how much data to sell at that price.
@@ V-D Benchmark Test
We conducted a benchmark test to compare the performance of our proposed quotation approach with the aforementioned baseline methods. Specifically, the quotation approach was configured to handle oversupply with the minor-sellers-first strategy. We tested all methods at different informed ratios from $0\%$ to $100\%$ with $10\%$ step length, and measured the server’s payoff, users’ payoff, and social welfare. The results of Monte Carlo tests are shown in Fig. 2.
@@
As we can see, our proposed quotation approach outperforms all the other methods in terms of social welfare. While the social welfare brought by the rigid GDPR solution drops at high informed ratio due to the high unlearning cost, both our quotation approach and the BSP solution exhibit an social welfare steadily increasing along with the informed ratio. Particularly, in comparison to BSP, our method significantly improves the server’s payoff at the cost of a less reduction in users’ payoff, leading to its superiority in the overall social welfare. This shall be explained in two aspects:
@@
    The multi-round quotation mechanism allows the server to explore the users’ non-linear privacy function over multiple intervals of data trading amount, which provides a closer approximation of the privacy utility loss by the payment. Compared to the BSP solution that provides only one optimum average price, the quotation mechanism is more cost-profitable for the buyer (server), but less profitable for the users.
@@ VI Conclusions
In this paper, we have proposed an iterative quotation-based price discovery framework for data redemption in machine unlearning. We have analyzed the incentives of buying and selling data of the server and users, respectively, and derived the optimal selling strategy of users. Compared with state-of-the-art baselines, our proposed quotation framework addresses a critical gap in deploying generative AI within next-generation mobile networks: balancing user data rights with operational imperatives. By enabling MNOs to dynamically compensate users for data retention, the mechanism sustains LLM accuracy–and thus network QoS–while respecting privacy preferences. Future work will explore: (1) taking into account users’ benefit in their selling strategy regarding network performance improvement; and (2) extending the model to scenarios with strategic information sharing among users regarding their data amount and privacy preferences.